OpenStudy (help!!!!):

. Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him? A. $496.30 B. $512.40 C. $555.40 D. $574.90 E. $588.20

OpenStudy (help!!!!):

steps are appreciated

OpenStudy (anonymous):

Let x be the original amount. To that you add 35% of x --> 0.35x The new amount is $670 Write an equation: x + (0.35)x = 670 (1 + 0.35)x = 670 1.35x = 670 x = 670/1.35 x ≈ 496.30 Answer: A. $496.30

OpenStudy (help!!!!):

how did you get 1 for x?

OpenStudy (help!!!!):

hey how did the person get 1 for x?

OpenStudy (anonymous):

1).Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him? Let original cost be "x" Then x+0.35x=670 1.35x=670 x= $496.30